1. In some patients, an allergic reaction may take the severe form of anaphylaxis. In such cases, epinephrine, a typically fast-acting bronchodilator, is sometimes administered by continuous IV infusion. If an adult dose is delivered at a rate of 4 micrograms per minute, and the concentration of the epinephrine solution is 1 mg/250 ml, then how many milliliters must be available for a one hour infusion?

2

Albuterol, a bronchodilator of the beta-2 adrenergic drug type, can be administered by nebulizer or inhaler. A nebulizer treatment for an adult patient consists of 2.5 mg of albuterol in the form of albuterol sulfate solution. This is administered from 5 to 15 minutes, three to four times a day. If the concentration of the solution is 1.25 mg/3 ml, how many milliliters are required for four administrations?
Answer #1

1.Given

In 1 minute=4mcg

So 60 minute =??mcg

Cross multiply and divide

=60×4÷1

=240mcg

So in 1 hour 240mcg will be infused

Now to calculate in ml

Available 1mg/250ml (1000mcg/250ml)

When 1000mcg=250ml

Then 240mcg=??ml

Cross multiply and divide

=240×250÷1000

=60000÷1000

=60

Ans:60ml will be available for 1 hour

2.Given

Medication:Albuterol

Order:2.5mg

Frequency:Four times a day

Time:5 to 15 minutes

Available:1.25mg/3ml

When 2.5mg=??ml

Cross multiply and divide

=2.5×3÷1.25

=7.5÷1.25

=6ml

In one dose 6ml to be administered

Therefore for four doses 6×4=24

Ans:24ml is required for four administration